# ISRE-$0$-$0$.00

### v2

@@
Recent advancements in generative AI, such as ChatGPT, have dramatically transformed how people access information. Despite its powerful capabilities, the benefits it provides may not be equally distributed among individuals—a phenomenon referred to as the digital divide. Building upon prior literature, we propose two forms of digital divide in the generative AI adoption process: (i) the learning divide, capturing individuals’ heterogeneous abilities to update their perceived utility of ChatGPT; and (ii) the utility divide, representing differences in individuals’ actual utility derived from per use of ChatGPT. To evaluate these two divides, we develop a Bayesian learning model that incorporates demographic heterogeneities in both the utility and signal functions. Leveraging a six-month clickstream dataset, we estimate the model and find significant learning and utility divides across various demographic attributes. Interestingly, lower-educated and non-white individuals derive higher utility gains from ChatGPT but learn about its utility at a slower rate. Furthermore, males, younger individuals, and those with an IT background not only derive higher utility per use from ChatGPT but also learn about its utility more rapidly. Besides, we document a phenomenon termed the belief trap, wherein users underestimate ChatGPT’s utility, opt not to use the tool, and thereby lack new experiences to update their perceptions, leading to continued underutilization. Our simulation further demonstrates that the learning divide can significantly affect the probability of falling into the belief trap, another form of the digital divide in adoption outcomes (i.e., outcome divide); however, offering training programs can alleviate the belief trap and mitigate the divide. Our findings contribute to the literature on the digital divide, AI adoption, and Bayesian learning. We also provide practical implications for governments, Non-profit organizations, AI providers, and companies aiming to promote equitable AI adoption.
@@ 1.1 Motivation
Search engines have long been the primary means by which individuals access online resources; over 93% of internet activities start with a search query [^1] See https://www.webfx.com/seo/statistics/. For instance, Google, which dominates the market with approximately 90% share, processes about 2 trillion searches annually [^2] See https://blog.hubspot.com/marketing/google-search-statistics. However, the recent emergence of large language models (LLMs) offers a new means to access these resources. Trained on extensive datasets (Zhang et al. 2024), applications such as ChatGPT can provide natural language responses to a wide range of user queries (Wang et al. 2023). The general-purpose nature of generative AI powered by LLMs (Morris et al. 2023), combined with their vast knowledge base, positions them as a compelling alternative to search engines (Xu et al. 2023).
Considering the significant potential of generative AI, one might anticipate widespread benefits. However, these potential advantages may be still constrained by the “digital divide,” which refers to the disparities in access, capabilities, and benefits derived from digital tools among individuals of varied backgrounds (Wei et al. 2011). Understanding and addressing this issue is crucial when aiming to transform technological progress into broader economic and societal advancements while ensuring equity and inclusiveness across the population (Burton-Jones and Sarker 2023, Simchi-Levi 2020).
Studying the digital divide is challenging due to both the complexity of the issue and the methodologies employed. Importantly, disparities in digital access and usage manifest differently across various technologies, making demographic differences context-dependent (Haslberger et al. 2023). As technology advances, research into these disparities has evolved from focusing on internet access to exploring various digital platforms and, more recently, artificial intelligence—leading to a new subfield termed the “AI divide” in the context of traditional AI (McElheran et al. 2024). However, given the uniqueness and power of LLM-based generative AI, which represents the first generation of artificial general intelligence (AGI) (Morris et al. 2023), it is imperative to understand the associated disparities and potentially initiate further research in the context of generative AI.
@@ 1.2 Research Questions and Findings
Building on previous research on the digital divide, we also conceptualize the GenAI divide in three layers (Wei et al. 2011, Dewan and Riggins 2005): (1) the access divide, referring to inequalities in technology access; (2) the capability divide, concerning individual differences in ability to effectively utilize the tools once accessed; and (3) the outcome divide, describing disparities in final outcomes, such as productivity and performance, which the preceding layers and other factors may influence. Furthermore, literature on IT and AI adoption has long emphasized the importance of individuals’ perceptions during the adoption process (e.g., perceived usefulness and perceived ease of use), as perceived benefits influence adoption decisions (Jo and Bang 2023, Venkatesh et al. 2003). However, perceptions are dynamic and inherently intangible, making it difficult to monitor how and why they change using data from surveys or experiments. Given these methodological challenges, it is understandable that previous digital divide research has not addressed disparities in individuals’ ability to effectively update their perceived utility through repeated interactions and other informational resources. We term this disparity the “learning divide,” a novel aspect of the second-layer capability divide. The learning divide is vital because varying abilities to update beliefs influence actual adoption and usage decisions, thereby affecting the welfare individuals ultimately receive (i.e., outcome divide). For example, if an individual erroneously believes the system is completely useless after a single trial, they may cease using it, thereby forfeiting opportunities to enhance their work and quality of life through the use of new technology. Conversely, if another person forms an accurate understanding after initial use, they can employ the system effectively and maximize their gains. To empirically investigate the existence of our proposed learning divide, we pose the following research question:
@@
RQ1 (learning divide): How do individuals with different demographic backgrounds update their perceived utility of generative AI at varying rates?
@@
Given that individuals’ perceptions are inherently unobservable and must be inferred from data, and their adoption decision-making is subject to informational uncertainty, structural learning models are ideally suited to address this problem (Wang and Yang 2024, Fang 2022). Therefore, to address the first research question, we develop a customized structural learning model to characterize individuals’ dynamic learning and adoption processes of generative AI, allowing for heterogeneous learning speeds related to their demographic attributes.
@@
In addition to the newly defined learning divide, individuals’ disparate utility derived from generative AI upon usage has become the focus of recent studies, although findings are mixed. On one hand, earlier research on AI’s impact on productivity suggests that workers with less prior expertise gain greater benefits from utilizing ChatGPT for specific tasks, potentially ameliorating existing inequalities (Noy and Zhang 2023). On the other hand, studies on AI adoption emphasize the significance of AI literacy (Ng et al. 2021), and note that performance disparities between older and younger groups have widened with the adoption of generative AI like ChatGPT (Haslberger et al. 2023). Most of these works are conducted in controlled experimental settings where participants are required to use specific tools to complete predetermined tasks. However, it remains underexplored how individuals perform in real-world contexts when they can potentially use general-purpose technologies for diverse tasks. To supplement previous studies, we aim to uncover the differences in utility gains when individuals from various backgrounds use generative AI tools. We term this the “utility divide,”, and pose the following research question:
@@
RQ2 (utility divide): How do individuals with different demographic backgrounds gain utility differently from generative AI when they opt to use it?
@@
Consistent with prior digital divide research employing behavioral experiments (Kohli et al. 2024, Haslberger et al. 2023, Robinson et al. 2020), our utility divide evaluates disparities in utility per use, conditional on adoption, constituting a specific type of third-level outcome divide. However, investigating this phenomenon in a real-world context is inherently challenging due to the natural feedback loop in the adoption process. When individuals believe a technology has high utility, they are more likely to adopt it. Upon using the generative AI tool, users update their beliefs based on their experiences (signals within the structural learning model), which in turn influence subsequent adoption decisions. Since individuals lack a priori knowledge of the actual utility, it influences their adoption decisions only indirectly through signal realizations, which are also governed by their learning abilities. Therefore, recognizing that the utility divide and learning divide are intertwined, we additionally incorporate utility heterogeneity into the structural learning framework to jointly model both disparities.
@@
To empirically estimate our model and address the abovementioned research questions, we obtained a dataset from a leading data collection firm. The dataset comprises online footprints of thousands of panelists over a six-month period following the launch of ChatGPT, enabling us to observe individual usage patterns over time. Additionally, the company provided demographic attributes for each participant, allowing us to explore heterogeneity. After fitting our model to the dataset, we uncovered significant and noteworthy learning and utility divides. Specifically, we find that male, younger, and individuals with an IT background both derive higher utility per use from ChatGPT and learn about its utility more rapidly. Besides, although individuals from minority racial groups or with lower educational attainment tend to derive greater utility gains per use, they unfortunately update their beliefs more slowly.
@@
Although we have empirically verified both the learning divide and the utility divide, it remains unclear how these disparities influence individuals’ lifetime usage and welfare dynamics—another form of outcome divides. The utility divide is, by definition, straightforward to interpret since it directly measures the utility gain per usage. However, the impact of the learning divide is less evident, as it requires translating differences in learning speed into variations in adoption dynamics. Furthermore, it remains unclear how effective counterfactual policies, such as training programs, might be in influencing these outcome divides through the learning process. Addressing this gap also aligns with the call for actionable insights in the digital divide literature (Wang et al. 2024, Wei et al. 2011). Therefore, we propose our final research question:
@@
To address this research question, we leverage our estimated model parameters to simulate individuals’ belief evolution trajectories over time. Through these simulations, we identify a phenomenon we term the “belief trap” wherein users who develop initially low perceptions of generative AI’s utility may cease usage, thereby forgoing opportunities to update their beliefs through further interactions. This self-reinforcing cycle bears striking similarities to the “poverty trap” in development economics, where individuals or nations become locked in persistent poverty without external intervention (Banerjee 2011). Our analysis reveals that users with disadvantaged learning abilities—as identified through our structural model—face significantly higher probabilities of falling into this belief trap, representing another manifestation of the outcome divide. Furthermore, our counterfactual simulations demonstrate that targeted training programs aimed at increasing exposure to generative AI can effectively mitigate disparities in trapping probabilities, indicating a viable policy intervention to address this form of the digital divide.
1.3 Contributions
This study makes several theoretical contributions. First, we present a unified framework for understanding IT adoption dynamics and the digital divide, particularly in scenarios characterized by information uncertainty and learning dynamics. This model is extensible to the adoption of other emerging technologies beyond generative AI. Second, we introduce two novel theoretical constructs—the learning divide and the belief trap—which are broadly applicable and offer new perspectives on technology adoption barriers. Third, within the emerging literature on generative AI, we are among the first to document demographic-based disparities in utility per use and learning abilities among individuals, quantifying how these differences manifest in adoption outcomes through our structural model.
Our paper also generates practical implications for multiple stakeholders. For generative AI providers, the presence of the utility gap underscores the imperative to implement inclusive design practices to enhance the tool’s value for underprivileged users. Moreover, our observation of learning disparity and belief trap suggests the importance of reducing information asymmetry through tutorials and training programs. This responsibility also extends to governments and non-profit organizations (NGOs). Those institutions can leverage our findings to design training programs and outreach initiatives that specifically target demographic groups we identified as susceptible to slower learning or belief traps.
@@ 1.4 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 offers a comprehensive literature review and situates our work within various research streams. Section 3 describes the data sources, variable construction, and exploratory analyses, underscoring the necessity of a structural learning framework. In Section 4, we present the model in detail, with estimation procedures and identification strategies outlined in Section 5. Section 6 provides the model estimation results and addresses RQ1 and RQ2. In Section 7, we perform further numerical experiments and evaluate counterfactual policy scenarios in response to RQ3. Finally, Section 8 concludes with broader discussions and highlights future research directions.
@@ 2 Related Literature
This research contributes to three streams of literature: the digital divide, AI adoption, and dynamic learning processes. In the subsequent subsections, we review findings from previous literature and highlight our study’s novel additions.
@@ 2.1 Digital Divide
The digital divide refers to disparities in access to, capability with, and effective use of information and communication technologies (ICT) among individuals from diverse backgrounds (Vassilakopoulou and Hustad 2023, OECD 2001). As a critical societal issue, it has been extensively researched (see reviews by Vassilakopoulou and Hustad (2023), Lythreatis et al. (2022), Scheerder et al. (2017), Riggins and Dewan (2005)), with key references summarized in Appendix Table LABEL:tab:digitaldivide. Owing to its complex, multidimensional nature, the digital divide manifests at three distinct levels (Wei et al. 2011, Dewan and Riggins 2005): (1) the first-level digital access divide, concerning inequalities in access to ICT; (2) the second-level digital capability divide, addressing disparities in the ability to effectively use ICT among those with access; and (3) the third-level digital outcome divide, describing differences in outcomes due to variations in access, usage patterns, and other related factors (Wei et al. 2011).
@@
The literature has identified a series of key factors contributing to the digital divide (Lythreatis et al. 2022, Van Dijk and Hacker 2003). These factors operate across multiple levels and dimensions, including socio-demographic and socio-economic characteristics (e.g., gender and age) (Venkatesh et al. 2014, Fairlie 2004, Jackson et al. 2001); personal attributes and skills (e.g., motivation and personality traits) (Celik 2023, Hsieh et al. 2011, Hargittai 2002); social and environmental influences (e.g., social support networks and geographical location) (Kohli et al. 2024, Hargittai 2003); and technical factors (e.g., platform affordance) (Wang et al. 2024, Yang and Zhang 2023). Among these, socio-demographic and socio-economic characteristics have consistently emerged as significant determinants of the digital divide (Venkatesh et al. 2014, Fairlie 2004, Jackson et al. 2001). Factors such as education level (Hidalgo et al. 2020, Hsieh et al. 2011, 2008), income (Fairlie 2004), age (Jackson et al. 2001), gender (Van Dijk and Hacker 2003, Hargittai 2003), and ethnicity (Fairlie 2004) have been shown to significantly influence individuals’ access to ICT, their ability to effectively use it, and the benefits they derive from ICT across different contexts and over time (Celik 2023, Venkatesh et al. 2014).
@@
However, it is important to note that the manifestation of the digital divide exhibit substantial variation across technologies and platforms (Yang and Zhang 2023, Burtch and Chan 2019), and that demographic effects on adoption and usage patterns may differ markedly depending on the specific technology or context being examined (Haslberger et al. 2023). As such, the digital divide is inherently a context-specific concept, which makes it necessary to specify the type of ICT under investigation (Vassilakopoulou and Hustad 2023). Consequently, the discourse surrounding the digital divide has evolved alongside the progression of digital technologies. Early studies focus on the unequal access to and use of personal computers (Autor et al. 1998), while later researcher expanded to include Internet access and use (Van Dijk and Hacker 2003, Hargittai 2002). Subsequently, the focus shifted towards the use of online services, such as e-government platforms (Venkatesh et al. 2014), and more recently has encompassed digital platforms including medical crowdfunding (Burtch and Chan 2019) and online rental housing markets (Boeing 2020).
The proliferation of AI technologies has led to a new dimension of digital inequality, the AI divide (McElheran et al. 2024). However, existing literature has primarily examined this divide within the context of narrow, task-specific traditional AI (Celik 2023), leaving a gap in our understanding of disparities in generative AI adoption. Generative AI such as ChatGPT, powered by large language models (LLMs), is distinguished by its foundation in general-purpose models and its ability to produce flexible, natural language outputs. This technology facilitates intuitive human-AI interactions that require minimal technical expertise from users, thereby broadening access for a diverse range of individuals. However, the outputs of generative AI are characterized by inherent variability (Morris et al. 2023). This uncertainty suggests that the utility derived from these systems can vary significantly among users with different backgrounds. Furthermore, effectively evaluating the utility of such AI systems may require an extensive learning process involving repeated interactions. Given the unique characteristics of generative AI, prior findings on the general digital divide or the specific AI divide may not be directly applicable to the context of generative AI.
Our study addresses this gap by developing a unified framework to address both the antecedents and the consequences of AI adoption. Because individuals’ perceived utility from AI influences their adoption decisions, and the realized outcomes subsequently affect their perceptions of AI, a natural feedback loop arises. However, due to the dynamic and intangible nature of perceptions, prior behavioral studies face methodological challenges in directly capturing the utility underlying participants’ adoption decisions or in addressing the dynamic adoption process. To move beyond the “snapshots” provided by survey studies, we develop a Bayesian learning model that offers theoretical underpinnings for the dynamics of the generative AI adoption process, complementing previous behavioral studies. Our model leverages a rich real-world dataset to recover the underlying drivers of adoption decisions (signals and perceived utility) and the direct consequences of individual adoption (actual utility). The model estimates can also be employed for counterfactual simulations to comprehensively evaluate indirect consequences, such as belief traps, and to assess policy interventions. Furthermore, our proposed model can be leveraged for future AI adoption studies involving other factors or technologies.

### v3

@@
Recent advancements in generative AI, such as ChatGPT, have dramatically transformed how people access information. Despite its powerful capabilities, the benefits it provides may not be equally distributed among individuals, a phenomenon referred to as the digital divide. Building upon prior literature, we propose two forms of digital divide in the generative AI adoption process: (i) the learning divide, capturing individuals’ heterogeneous abilities to update their perceived utility of ChatGPT; and (ii) the utility divide, representing differences in individuals’ actual utility derived from per use of ChatGPT. To evaluate these two divides, we develop a Bayesian learning model that incorporates heterogeneities in both the utility and signal functions. Leveraging a large-scale clickstream dataset, we estimate the model and find significant learning and utility divides across various social characteristics. Interestingly, individuals without any college education, non-white individuals, and those with lower English literacy derive larger utility gains from ChatGPT, yet update their beliefs about its utility at a slower rate. Furthermore, males, younger individuals, and those in occupations with greater exposure to generative AI not only obtain higher utility per use from ChatGPT but also learn about its utility more rapidly. Besides, we document a phenomenon termed the belief trap, wherein users underestimate ChatGPT’s utility, opt not to use the tool, and thereby lack new experiences to update their perceptions, leading to continued underutilization. Our simulation further demonstrates that the learning divide can significantly affect the probability of falling into the belief trap, another form of the digital divide in adoption outcomes (i.e., outcome divide); however, offering training programs can alleviate the belief trap and mitigate the divide. Our findings contribute to the literature on the digital divide, AI adoption, and Bayesian learning.
@@ 1.1 Motivation
Search engines have long been the primary means by which individuals access information; over 93% of internet activities start with a search query [^1] See https://www.webfx.com/seo/statistics/. For instance, Google, which dominates the market with approximately 90% share, processes about 2 trillion searches each year [^2] See https://blog.hubspot.com/marketing/google-search-statistics. However, the recent emergence of large language models (LLMs) offers a new means to access these resources. Trained on extensive datasets (Zhang et al. 2024), LLM-backboned generative AI, such as ChatGPT, can provide natural language responses to a wide range of user queries (Wang et al. 2023). The general-purpose nature of generative AI powered by LLMs (Morris et al. 2023), combined with their vast knowledge base, positions them as a compelling alternative to search engines (Xu et al. 2023).
Given the significant potential of generative AI, one might anticipate widespread benefits. However, these potential advantages may still be constrained by the “digital divide,” which refers to the disparities in access to, capability with, and effective use of information and communication technologies (ICT) among individuals from diverse backgrounds (Vassilakopoulou and Hustad 2023, OECD 2001). Understanding and addressing the digital divide is crucial when aiming to transform technological progress into broader economic and societal advancements across the population (Burton-Jones and Sarker 2023, Simchi-Levi 2020).
However, the digital divide is often context dependent and can manifest differently across technologies (Haslberger et al. 2023). As technologies have advanced, research into digital disparities has expanded from internet access to various digital platforms and, more recently, artificial intelligence (AI), giving rise to a new subfield termed the “AI divide” (McElheran et al. 2024). Compared to earlier ICTs, LLM-based generative AI exhibits distinctive properties, raising doubts about whether findings from prior ICT contexts can apply. First, its natural-language interface enables highly customizable inputs and produces outputs that are relatively easy to consume, which may help address users’ information needs efficiently, particularly in long-tail domains (Xu et al. 2023). At the same time, system outputs are typically unstructured, making the assessment of model capabilities more salient and, in most settings, more difficult (Zhong et al. 2024). Moreover, system outputs may fail to align with users’ intentions and exhibit hallucinations, further increasing noise in delivered information (Xu et al. 2024). Given these unique properties, it is essential to examine how they translate into downstream disparities in adoption and, more broadly, to extend digital divide research to the context of generative AI (thereafter generative AI divide).
@@ 1.2 Research Questions and Findings
Building on previous research on the digital divide, we also conceptualize the generative AI divide in three layers (Wei et al. 2011, Dewan and Riggins 2005): (1) the access divide, referring to inequalities in access to generative AI; (2) the capability divide, concerning individual differences in ability to utilize generative AI once accessed; and (3) the outcome divide, describing disparities in final outcomes, such as productivity and performance, which the preceding layers and other factors may influence. Furthermore, literature on IT and AI adoption has long emphasized the importance of individuals’ perceptions during the adoption process (e.g., perceived usefulness and perceived ease of use), as perceived benefits influence adoption decisions (Jo and Bang 2023, Venkatesh et al. 2003). However, perceptions are dynamic and intangible, making it difficult to monitor how and why they change using data from surveys or experiments. Such dynamics are particularly salient in the generative AI context, where evaluating system capability itself is challenging and constantly evolving (Zhong et al. 2024) and where alignment related issues introduce additional noise (Xu et al. 2024).
Therefore, individuals’ ability to update their perceptions of generative AI can substantially shape their adoption process. This view is consistent with the literature on AI literacy, which refers to the ability to understand, use, and evaluate AI systems (Annapureddy et al. 2025, Ng et al. 2021). Because AI literacy varies widely across individuals with different social characteristics (Wang et al. 2024, Celik 2023, Robinson and others 2020), we aim to quantitatively measure and document disparities in individuals’ ability to update their perceived utility through repeated interactions and other informational resources. We term this disparity the “learning divide,” a novel aspect of the second layer capability divide. The learning divide is consequential because heterogeneity in belief updating shapes adoption and usage decisions, thereby affecting the welfare individuals ultimately receive (i.e., the outcome divide). For example, if an individual mistakenly concludes that the system is entirely useless after a single trial, they may stop using it and forgo opportunities to improve their work and quality of life through the technology. Conversely, if another individual forms an accurate understanding after initial use, they can use the system effectively and realize greater gains. To empirically examine the learning divide, we pose the following research question:
@@
RQ1 (learning divide): How do individuals with different social characteristics update their perceived utility of generative AI at varying rates?
@@
Given that individuals’ perceptions are unobservable and must be inferred from data, and their adoption decision-making is subject to informational uncertainty, structural learning models are ideally suited to address this problem (Wang and Yang 2024, Fang 2022). Therefore, to address the first research question, we develop a customized structural learning model to characterize individuals’ dynamic learning and adoption processes of generative AI, allowing for heterogeneous learning speeds related to their social characteristics.
@@
In parallel, individuals’ disparate utility derived from generative AI upon usage has become the spotlight of recent studies in generative AI adoption, but now findings are mixed. On the one hand, earlier research on AI’s impact on productivity suggests that workers with less prior expertise gain greater benefits from utilizing ChatGPT for specific tasks, potentially ameliorating existing inequalities (Noy and Zhang 2023). On the other hand, studies on AI adoption note that performance disparities have widened with the adoption of generative AI like ChatGPT (McElheran et al. 2024, Haslberger et al. 2023). One of the potential explanations for the mixed findings is that most of these works are conducted in controlled experimental settings where participants are required to use one of specific generative AI tools to complete predetermined tasks. However, it remains underexplored how individuals perform in real-world settings where they can use those general-purpose technologies for diverse tasks. To supplement previous studies, we aim to uncover the differences in utility gains when individuals with various social characteristics use generative AI tools. We term this the “utility divide,” and pose the following research question:
@@
RQ2 (utility divide): How do individuals with different social characteristics gain utility differently from generative AI when they opt to use it?
@@
Consistent with research on digital divide employing behavioral experiments (Kohli et al. 2024, Haslberger et al. 2023, Robinson and others 2020), our utility divide evaluates disparities in utility per use, conditional on adoption, constituting a specific type of third-level outcome divide. However, investigating this phenomenon in a real-world context is challenging due to the natural feedback loop in the adoption process. When individuals believe that a technology has high utility, they are more likely to adopt it. Upon using the generative AI tool, users update their beliefs based on their experiences (signals within the structural learning model), which in turn influence subsequent adoption decisions. Since individuals lack a priori knowledge of the actual utility, it influences their adoption decisions only indirectly through signal realizations, which are also governed by their learning abilities. Therefore, recognizing that the utility divide and learning divide are intertwined, we incorporate utility heterogeneity into the structural learning framework to jointly model both disparities.
@@
To empirically estimate our model and address the abovementioned research questions, we obtain a dataset from a leading data collection firm. The dataset comprises online footprints of thousands of panelists over a six-month period following the launch of ChatGPT, enabling us to observe individual usage patterns over time. Additionally, the company also provides social characteristics for each participant, allowing us to explore how social characteristics shape the proposed divides. After fitting our model to the dataset, we uncover significant learning and utility divides. Notably, individuals without college education, non-white individuals, and those with lower English literacy realize larger utility gains from ChatGPT, yet revise their beliefs about its utility more slowly. In contrast, males, younger individuals, and workers in occupations with greater exposure to LLMs both extract higher utility per use and update their perceived utility more quickly.
@@
Although we have empirically validated both the learning divide and the utility divide, it remains unclear how these disparities influence individuals’ lifetime usage and welfare dynamics, another form of outcome divides. The utility divide is, by definition, straightforward to interpret since it directly measures the utility gain per usage. However, the impact of the learning divide is less evident, as it requires translating differences in learning speed into variations in adoption dynamics. Furthermore, it remains unclear how effective counterfactual policies might be in influencing these outcome divides through the learning process. Addressing this gap also aligns with the call for actionable insights in the literature on digital divide (Wang et al. 2024, Wei et al. 2011). Therefore, we propose our final research question:
@@
To address this research question, we leverage our estimated model parameters to simulate individuals’ belief evolution trajectories over time. Through these simulations, we identify a phenomenon we term the “belief trap”, in which users who initially form low perceptions of generative AI’s utility may cease using it, thereby forgoing opportunities to update their beliefs through further interactions. This self-reinforcing cycle bears striking similarities to the “poverty trap” in development economics, where individuals or nations become locked in persistent poverty without external intervention (Banerjee 2011). Our analysis further shows that users with disadvantaged learning ability, as identified by our structural model, are substantially more likely to fall into this belief trap, constituting another manifestation of the outcome divide. We also quantify the extent of news exposure or product improvement required to reduce belief trap risk and sustain long-run adoption, underscoring how difficult it can be to narrow these divides. Finally, our counterfactual simulations indicate that targeted training programs that increase exposure to generative AI can meaningfully lower trapping probabilities and reduce disparities, suggesting a feasible policy intervention to address this new form of digital divide.
1.3 Theoretical Contributions
This study makes several theoretical contributions. First, we present a unified framework for understanding IT adoption dynamics and the digital divide, particularly in scenarios characterized by information uncertainty and learning dynamics. Second, to empirically test this framework, our study leverages large-scale real-world usage data in which interaction with generative AI is voluntary, repeated, and consequential. This setting allows us to observe how beliefs about system capability evolve endogenously over time, how early noisy experiences shape subsequent adoption decisions, and how users self-select into or out of learning opportunities. As a result, we are able to uncover dynamic mechanisms that are mechanically ruled out in experimental designs with mandated exposure. Therefore, our paper presents a complementary approach to prior experimental studies that examine short-run productivity or learning effects under controlled exposure (Humlum and Vestergaard 2024, Noy and Zhang 2023, Haslberger et al. 2023). Third, our paper introduces two novel theoretical constructs, the learning divide and the belief trap, which are broadly applicable and offer new perspectives on technology adoption barriers. Fourth, with regard to the emerging literature on generative AI, we are among the first to document social characteristics-based disparities in both utility per use and learning ability and to quantify how these differences translate into adoption outcomes through a structural model. Finally, we map our empirical estimates into counterfactual policy evaluations, contributing evidence on the effectiveness of potential interventions to mitigate divides in generative AI adoption.
@@ 1.4 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 reviews the related literature and positions our study within past research streams. Section 3 then develops the theoretical arguments about the generative AI divide. Section 4 describes the data sources, variable construction, and exploratory analyses. Section 5 presents the model, while Section 6 details the estimation procedures and identification strategy. Section 7 reports the estimation results and addresses RQ1 and RQ2. In Section 8, we conduct numerical experiments and evaluate counterfactual scenarios to address RQ3. Finally, Section 9 concludes with a broader discussion and outlines directions for future research.
@@ 2 Related Literature
Our paper contributes to three streams of literature: digital divide, AI adoption, and dynamic learning processes. In each subsection, we review findings from previous literature, identify the research gap, and highlight how our study extends the literature.
@@ 2.1 Digital Divide
The digital divide refers to the disparities in access to, capability with, and effective use of information and communication technologies (ICT) among individuals from diverse backgrounds (Vassilakopoulou and Hustad 2023, OECD 2001). As a critical societal issue, it has been extensively investigated (see reviews by Vassilakopoulou and Hustad (2023), Lythreatis et al. (2022), Scheerder et al. (2017), Dewan and Riggins (2005)), with key references summarized in Appendix Table LABEL:tab:digitaldivide. Reflecting its complex, multidimensional nature, the digital divide manifests at three distinct levels (Wei et al. 2011, Dewan and Riggins 2005): (1) the first-level access divide, concerning inequalities in access to ICT; (2) the second-level capability divide, addressing disparities in the ability to effectively use ICT among those with access; and (3) the third-level outcome divide, describing differences in outcomes due to variations in access, usage patterns, and other related factors (Wei et al. 2011).
@@
Prior studies have revealed a series of key factors contributing to the digital divide across multiple levels and dimensions, including social characteristics (Venkatesh et al. 2014, Fairlie 2004, Jackson et al. 2001); personal attributes and skills (e.g., motivation and personality traits) (Celik 2023, Hsieh et al. 2011, Hargittai 2002); social and environmental influences (e.g., social support networks and geographical location) (Kohli et al. 2024, Hargittai 2003); and technical factors (e.g., platform affordance) (Wang et al. 2024, Yang and Zhang 2023). Among them, social characteristics have been identified as central determinants of the digital divide (Venkatesh et al. 2014, Fairlie 2004, Jackson et al. 2001). The social characteristics, including education (Hidalgo et al. 2020, Hsieh et al. 2008), age (Jackson et al. 2001), gender (Van Dijk and Hacker 2003, Hargittai 2003), race (Fairlie 2004), language proficiency (Bassignana et al. 2025, Hargittai 2003, 2002), and occupational position and job roles (McElheran et al. 2024, Venkatesh et al. 2014), have been shown to shape individuals’ access to ICT, their ability to use it effectively, and the outcomes they derive across contexts and over time (Celik 2023, Scheerder et al. 2017). Importantly, prior literature emphasizes digital literacy as a key mechanism through which social characteristics translate into differential ICT usage and outcomes, mediating the relationship between social characteristics and realized benefits (Scheerder et al. 2017, Hargittai 2002). This mechanism is especially salient in the context of generative AI, where effective use depends on formulating appropriate prompts, interpreting probabilistic and unstructured outputs, and evaluating system capability under uncertainty.
@@
In addition, the manifestation of the digital divide exhibits substantial variation across technologies and platforms (Yang and Zhang 2023, Burtch and Chan 2019), and the effects of social characteristics on adoption and usage patterns may differ depending on the specific technology or context (Haslberger et al. 2023). As such, the digital divide is a context-specific concept, which makes it necessary to specify the type of ICT under investigation (Vassilakopoulou and Hustad 2023). As a result, the discussion about the digital divide has evolved alongside the progression of digital technologies. Early studies focus on the unequal access to and use of personal computers (Autor et al. 1998), while later researchers expanded to include Internet access and use (Van Dijk and Hacker 2003, Hargittai 2002). Then, the focus shifted towards the use of digital platforms, such as e-government platforms (Venkatesh et al. 2014), medical crowdfunding platforms (Burtch and Chan 2019), and online rental housing platforms (Boeing 2020). Most recently, the diffusion of AI has given rise to a new dimension of inequality, the AI divide (McElheran et al. 2024). Understanding this divide in the context of generative AI is particularly urgent, as it dramatically differs from earlier ICTs in its probabilistic, adaptive, and interactive nature. Therefore, it is meaningful to explore what new forms of inequality these distinctive technical properties generate and how social characteristics shape capability and outcome divides in the context of generative AI.